# TaskEval: Assessing Difficulty of Code Generation Tasks for Large Language Models

### v2

@@ 4.1. Hyperparameters used
To have different prompts, we choose to generate 3 levels and 6 rephrasings per level. We chose those parameters as the trade-off between generating diverse enough prompts while limiting the amount to generate. For a benchmark of $n$ tasks, this effectively means we have to generate $18\times n$ prompts. We further ask each CodeLLM to generate 5 code samples per individual prompt resulting in $90\times n$ code samples for each CodeLLM to generate. As an example, for a benchmark containing 200 tasks, this process yields 18,000 generated code for our evaluation per CodeLLM. We make available all the prompts and code generated in our replication package (rep, 2024).
@@ 4.2. Benchmark used
In our experiment, we chose two benchmarks: HumanEval+ (Liu et al., 2023) and ClassEval (Du et al., 2023). Those benchmarks were chosen as they cover different types of tasks and dependency levels. HumanEval+ is widely used as a standard benchmark in code generation (Big, 2024a). It is an extended variant of HumanEval (Chen et al., 2021), that contains more tests to assess the correctness of code samples. The benchmark consists of 164 handwritten Python programming problems with a function signature, docstring, body and several unit tests. HumanEval+ tasks are at the function level dependency as they do not require external libraries, besides Python standard library, or class/file contexts to be operated. For the original prompt, we use the docstring as given in HumanEval+.
ClassEval is a benchmark of 100 handcrafted classes coded in Python. Each class contains a description and on average 4 methods each with a docstring. We consider a task to be one of the 400 methods of ClassEval. Contrary to HumanEval+, ClassEval tasks are more similar to practical code used in projects and rely on a class context to operate. ClassEval proposes three prompting strategies: holistic, incremental and compositional. In our case, we use the compositional strategy, that is, providing the class context to the LLM and asking it to complete one method of the class, only modifying the docstring of the method to complete through our crafted prompts. We do not make use of the other generation approaches as it reduces the control we have over the prompts and could introduce biases: the holistic approach, as it implies asking the LLM to generate the whole class from scratch, can impact the performance artificially. Similarly, the incremental approach requires the reuse of codes generated by the LLM for other methods when generating new ones, which could also alter the difficulty. As we need to generate a high number of code samples and prompts, we reduce the number of tasks to evaluate by doing a stratified sampling at the 95% confidence interval. For 400 methods, this gives us 197 methods to sample which we round up to 200. Those 200 methods are sampled equally across classes resulting in 2 methods per class available. Methods chosen are available in our replication package (rep, 2024).
@@ 4.3. Models used
In our experiments, we use GPT-4 (GPT, 2023b) as both the Prompt LLM (to generate the different prompts) and the Task LLM (to generate new hard tasks). We do so as GPT-4 is the state-of-the-art LLM and so should provide answers with the best quality which would require a minimum human supervision. For the CodeLLM, we use 5 LLMs with different architecture: CodeLLama 7B (Roziere et al., 2023), MagiCoder 6.7B (Wei et al., 2023), DeepSeekCode 7B (Guo et al., 2024), CodeGemma 7B (Team et al., 2024) and GPT-3.5 (GPT, 2023a). We have not used GPT-4 as CodeLLM to avoid biases as it is used to generate the prompts fed to the CodeLLM. For all models, we used their instructions-tuned versions so that the models can handle the different prompts properly. This gives us a diverse array of architecture and training procedures. For all code generation in TaskEval, since we are using sampling generation, we set the temperature hyperparameter to 0.8 as it is one of the temperature settings commonly used in other similar studies (Liu et al., 2024a; Chen et al., 2021; Liu et al., 2023; Roziere et al., 2023). In all cases except GPT models (for which we use OpenAI API), we use the HuggingFace (HFm, 2024) implementation of the models to allow for replication. Ranking of the LLMs’ performance follows the assessment obtained on the original benchmarks.
@@ 4.4. IRT model fit
For the IRT model, we use the library Birt-GD (Ferreira-Junior et al., 2023) which contains a gradient descent based Maximum Likelihood Estimation of the $\beta^{3}$-IRT algorithm. We use the implementation exactly as is, except we modify the initial values of the parameters: $\delta_{j}$ and $\theta_{i}$ are initialized as the average of scores for each task/CodeLLMs while $a_{j}$ is initialized as the Pearson correlation coefficient between the $\theta_{i}$ and the scores on each task of the $i^{th}$ CodeLLM. This last initialization helps to solve the symmetry issue of the $\beta^{3}$-IRT model (Ferreira-Junior et al., 2023). The implementation details can be found in our replication package (rep, 2024). We use the scores collected as described in Section 2. Then, we will fit two IRT models, one for HumanEval and one for ClassEval data. The $R^{2}$ (Ferreira-Junior et al., 2023) after fitting the IRT models to our data are $0.928$ and $0.927$, respectively, showing an accurate fit.
@@ 4.5. Research Questions
RQ1: How do tasks’ characteristics varies across benchmark? In our first RQ, we will analyze the tasks of both benchmarks through the prism of our IRT model. To do so, we will first compare the two benchmarks based on the expected probability of response of each CodeLLM, that is $E[p_{ij}|\theta_{i},\delta_{j},a_{j}]$, through their cumulative distributions across tasks. Besides observations of the cumulative distributions, we apply the Anderson-Darling (AD) (Scholz and Stephens, 1987) test, a statistical test to assess the difference between the two empirical distributions, between the distributions across benchmarks for the same CodeLLM. Then, using the obtained discriminant $a_{j}$ and difficulty $\delta_{j}$, as calculated by the IRT model, we will draw a portrait, task-wise, of both benchmarks. This will allow us to describe tasks present in the benchmark qualitatively.
@@
RQ2: Do tasks’ topics affect their characteristics? In this RQ, we explore how tasks’ topics can affect how the difficulty/discriminant parameters are evaluated. We chose to explore at the topic level as it gives a good trade-off between a more nuanced analysis than the benchmark level while allowing us, compared to individual tasks analysis, to draw some conclusions over a group of tasks. To obtain meaningful topics, we will leverage the topic modeling approach with BERTopic (Grootendorst, 2022), as was used in similar studies when dealing with natural language data (Hämäläinen et al., 2023; Fütterer et al., 2023; Shen et al., 2023). For each benchmark, the topic modeling clusters all tasks using the prompts of Level 1 before rephrasing. We use this specific prompt type to have similar formatting across tasks while being as close as possible to the original prompt as mentioned in Section 3.2. We consider a topic to be valid only if it contains at least three tasks. This criterion ensures that in ClassEval, we avoid ending up with an excessive number of topics, each containing only two tasks for example, which could occur due to the sampling methodology used to select tasks from the benchmark. As the process will inherently label some tasks as “noise”, noisy tasks are not considered in the analysis. This results in 17 topics for HumanEval+ and 21 topics for ClassEval. For the topics’ names, we use the names provided by BERTopic as a base before having the first two authors refine them by checking the tasks in the topic. Those topics constitute tasks that have a similar functionality/goal, for instance, tasks dealing with “sequence generation” or “SQL statements”. Then, for each topic, the mean accuracy of each CodeLLM as well as the difficulty and discriminant among all tasks in the topic is calculated and analyzed.
@@
RQ3: Do programming constructs affect tasks’ characteristics? In this RQ, we aim to explore the relationship between the difficulty or discriminative power of a task for a CodeLLM and the program constructs (e.g., If statements, For-loops, Function Calls) present in the generated code snippets used to solve the task. To achieve this, we will analyze the Abstract Syntax Tree (AST) of each code fragment generated for the task. By parsing these fragments, we will compute the average frequency of each program construct per task, enabling us to identify potential correlations between specific constructs and task complexity. To ensure that the selected tasks are not dominated by irrelevant or incorrect code fragments, we retain only those tasks for which, across all CodeLLMs, at least half of the generated samples are similar to the correct solution. This threshold of 50% similarity ensures that the majority of the code fragments are aligned with the intended task, even if they were generated from a single type of prompt. While we could have restricted our selection exclusively to code fragments that pass all tests (i.e., fully correct solutions), we opted to include similar-to-correct snippets to preserve a broader diversity of constructs. Indeed, code samples can be wrong to different degrees depending on the bugs affecting the code sample (Tambon et al., 2024). It could also bias the analysis towards code fragments obtained for higher-level prompts that are more likely to generate correct code. However, we also should not take any code fragments, as we may end up with generated code not even trying to solve the task if the CodeLLM misunderstood the prompt. To strike a balance, we will only consider code fragments that are similar enough to a correct one. We do so by using CodeBLEU (Ren et al., 2020) which measures the syntactical correctness of code samples. We thus measure the similarity between each code fragment and a correct code fragment that passed the tests. Note that, there is always one correct code fragment per task as a reference which is provided in the benchmark. We chose CodeBLEU as it is a widely used metric in code generation tasks (Lu et al., 2021; Mastropaolo et al., 2023; Lozhkov et al., 2024; Wang et al., 2021). Regarding the decision threshold, in the original paper, CodeBLEU’s authors found a correlation between human similarity assessment and CodeBLEU similarity on text-to-code generation, with a value of $0.3$ for CodeBLEU representing an average subjective human assessment of $3$ out of $5$ (Ren et al., 2020). In our case, we take $0.5$ as a threshold to select/reject code fragments. This helps us preserve code snippets that are false but still address the task while discarding junk code fragments.
This leaves us with $81\%$ of HumanEval+ tasks and $86\%$ of ClassEval tasks that are usable for the analysis. Though tasks cut out of the analysis are mostly in very high difficulty ($>0.9$), we still have multiple tasks with a high difficulty $>0.5$ for the analysis. Finally, we extract the AST nodes in each of the code fragments of the tasks considered using the ast package from the Python standard library. For each node extracted in that way, we register the node type as per the AST. Once the nodes are extracted, as the number of lines of code could bias the results (e.g., if harder tasks require more lines of code to be solved), we normalize by the number of lines of code of each fragment. Finally, we calculate the average number of each program construct across all code fragments within the same task. This gives us the number of each node type, across tasks and CodeLLMs, which will be compared with the difficulty/discriminant of each task using the Kendall-$\tau$ (Kendall, 1938) to measure the correlation while handling ties. The node types with a positive significant correlation will signify an increasing trend with difficulty. Furthermore, we will also use the AD test to check whether the distributions of the top/bottom 50% difficult/discriminant tasks exhibit different behavior for each programming construct and each CodeLLMs.
@@
RQ4: Do human annotators and CodeLLMs rate tasks’ difficulty similarly? In our final RQ, we aim to compare task difficulty assessments made by human annotators with those generated by our approach. To achieve this, we collect human-labeled difficulty ratings for both HumanEval+ and ClassEval by employing paid developers as crowd workers on the Prolific platform [^1] https://www.prolific.com/, which was used in previous studies (Zid et al., 2024). To have the crowd-workers estimate the difficulty, the most straightforward approach would be to have them implement the selected tasks. However, a pilot study with 10 developers for 12 tasks we conducted revealed that several participants are inclined to use LLMs to code for themselves despite it being forbidden as advertised in the study. Given our limited control over how recruited participants might use LLMs, we adopted an alternative experimental setup that leverages effort estimation techniques, such as Planning Poker (Grenning, 2002), which is familiar to developers. Since our study involves a crowd of workers, we followed the experimental design proposed by Alhamed et al. (Alhamed and Storer, 2021), which focuses on effort estimation when relying on a crowd of non-expert. Their approach simulates real-world effort estimation processes but eliminates the need for developers to resolve estimation conflicts. The method is structured into iterative rounds. In the initial round, individual participants are asked to estimate the difficulty of a task through a time-based scale (e.g., one hour or one day) using solely the information included in the task description. Participants are also required to provide comments justifying their estimates. Once a sufficient number of responses are collected, the level of agreement among participants is calculated. If the agreement exceeds a predefined threshold, the task’s effort estimation is determined by taking the median of the responses. If the threshold is not met, a subsequent round is conducted with a new set of participants. These participants are provided with the task information, along with the comments and results from the previous round. This process repeats until either an agreement is reached or a maximum number of rounds is completed.
@@
In this study, we implemented Alhamed et al.’s approach with $5$ participants per task for a given round and a threshold of agreement level set to $0.6$ (i.e., Substantial Agreement). Given that we have a limited budget (as crowd-workers are paid based on the time spent), we aimed for $50$ developers in the first round. This results in $120$ tasks (out of the $364$ of both benchmarks) being considered, which were sampled at random. Tasks were then sorted into $10$ questionnaires of $12$ assessments ($6$ for HumanEval+ and $6$ for ClassEval). To balance the time required to complete the questionnaires, we categorized tasks into three groups based on the number of lines of code and the cyclomatic complexity of their oracles before sampling. All questionnaires can be found in our replication package (rep, 2024). Each assessment is decomposed into several parts: First, several prompts for the same tasks are provided to the participants, which have to summarize what the task is about. This allows us to check that the participant properly read the task. Then, the participant is invited to estimate the duration it would take them (from 1 minute to over 40 minutes) as well as the subjective difficulty of the task (using a Likert scale from ”Very Easy” to ”Very Difficult”). They then had to explain their choice, based on considerations such as the time it would take to look for a particular concept, and how to use a given library. At the beginning of the questionnaire, participants were given two examples to clarify expectations. Responses were collected until $5$ participants provided acceptable answers. To ensure quality, we manually reviewed responses, verifying that participants provided meaningful explanations for at least two items (e.g., avoiding one-word answers for reasoning questions or random difficulty/time estimations). In particular, while we rely on the time estimation for our analysis but still ask for the subjective difficulty, we could analyze the trend between the two variables for each participant and potentially detect incoherence (e.g., Easy tasks with 40 minutes estimation and Hard tasks with 10 minutes) which could signal random filling of the questionnaire. We followed Alhamed et al. (Alhamed and Storer, 2021) procedure until all tasks have been assessed with a substantial agreement. We then analyzed the difference between human-assessed difficulty and the difficulty for CodeLLMs obtained using our IRT-based approach. Since human assessments are discrete, we discretized the difficulty scores obtained via IRT within the interval [0,1] for comparison.

### v3

@@ 4.1. Hyperparameters used
To have different prompts, we choose to generate 3 levels and 6 rephrasings per level. We chose those parameters as a trade-off between generating diverse enough prompts while limiting the number to generate. For a benchmark of $n$ tasks, this effectively means we have to generate $18\times n$ prompts. We further ask each CodeLLM to generate 5 code samples per individual prompt, resulting in $90\times n$ code samples for each CodeLLM to generate. As an example, for a benchmark containing 200 tasks, this process yields 18,000 generated code for our evaluation per CodeLLM. We make available all the prompts and code generated in our replication package (rep, 2024).
@@ 4.2. Benchmark used
In our experiment, we chose two benchmarks: HumanEval+ (Liu et al., 2023) and ClassEval (Du et al., 2023). Those benchmarks were chosen as they cover different types of tasks and dependency levels. HumanEval+ is widely used as a standard benchmark in code generation (Big, 2024a). It is an extended variant of HumanEval (Chen et al., 2021) that contains more tests to assess the correctness of code samples. The benchmark consists of 164 handwritten Python programming problems with a function signature, docstring, body and several unit tests. HumanEval+ tasks are at the function level dependency as they do not require external libraries, besides the Python standard library, or class/file contexts to operate. For the original prompt, we use the docstring as given in HumanEval+.
ClassEval is a benchmark of 100 handcrafted classes coded in Python. Each class contains a description and, on average, 4 methods each with a docstring. We consider a task to be one of the 400 methods of ClassEval. Contrary to HumanEval+, ClassEval tasks are more similar to practical code used in projects and rely on a class context to operate. ClassEval proposes three prompting strategies: holistic, incremental and compositional. In our case, we use the compositional strategy, that is, providing the class context to the LLM and asking it to complete one method of the class, only modifying the docstring of the method to complete through our crafted prompts. We do not make use of the other generation approaches as it reduces the control we have over the prompts and could introduce biases: the holistic approach, as it implies asking the LLM to generate the whole class from scratch, can impact the performance artificially. Similarly, the incremental approach requires the reuse of codes generated by the LLM for other methods when generating new ones, which could also alter the difficulty. As we need to generate a high number of code samples and prompts, we reduce the number of tasks to evaluate by doing a stratified sampling at the 95% confidence interval. For 400 methods, this gives us 197 methods to sample, which we round up to 200. Those 200 methods are sampled equally across classes, resulting in 2 methods per class available. Methods chosen are available in our replication package (rep, 2024).
@@ 4.3. Models used
In our experiments, we use GPT-4 (GPT, 2023b) as the Prompt LLM (to generate the different prompts). We do so as GPT-4 is the state-of-the-art LLM111at the time the experiments were started and so should provide answers with the best quality, which would require a minimum of human supervision. For the CodeLLM, we use 8 LLMs with different architectures: YiCoder-1.5B (AI et al., 2025), CodeLLama 7B (Roziere et al., 2023), MagiCoder 6.7B (Wei et al., 2023), DeepSeekCoder 7B (Guo et al., 2024), CodeGemma 7B (Team et al., 2024), QwenCoder-2.5 7B (Yang et al., 2024), GPT-3.5 (GPT, 2023a), DeepSeek-V3 (DeepSeek-AI, 2024). We have not used GPT-4 as CodeLLM to avoid biases, as it is used to generate the prompts fed to the CodeLLM. For all models, we used their instructions-tuned versions so that the models can handle the different prompts properly. This gives us a diverse array of architecture and training procedures. For all code generation in TaskEval, since we are using sampling generation, we set the temperature hyperparameter to 0.8 as it is one of the temperature settings commonly used in other similar studies (Liu et al., 2024a; Chen et al., 2021; Liu et al., 2023; Roziere et al., 2023). In all cases except GPT models (for which we use OpenAI API) and DeepSeek-V3 (for which we use DeepSeek API), we use the HuggingFace (HFm, 2024) implementation of the models to allow for replication. Ranking of the LLMs’ performance on our prompts follows the assessment obtained on the original benchmarks.
@@ 4.4. IRT model fit
For the IRT model, we use the library Birt-GD (Ferreira-Junior et al., 2023) which contains a gradient descent-based Maximum Likelihood Estimation of the $\beta^{3}$-IRT algorithm. We use the implementation exactly as is, except we modify the initial values of the parameters: $\delta_{j}$ and $\theta_{i}$ are initialized as the average of scores for each task/CodeLLMs while $a_{j}$ is initialized as the Pearson correlation coefficient between the $\theta_{i}$ and the scores on each task of the $i^{th}$ CodeLLM. This last initialisation helps to solve the symmetry issue of the $\beta^{3}$-IRT model (Ferreira-Junior et al., 2023). The implementation details can be found in our replication package (rep, 2024). We use the scores collected as described in Section 2. Then, we will fit two IRT models, one for HumanEval and one for ClassEval data. The $R^{2}$ (Ferreira-Junior et al., 2023) after fitting the IRT models to our data are $0.87$ and $0.90$, respectively, showing an accurate fit.
@@ 4.5. Research Questions
RQ1: How do tasks’ characteristics varies across benchmark? In our first RQ, we will analyse the tasks of both benchmarks through the prism of our IRT model. To do so, we will first compare the two benchmarks based on the expected probability of response of each CodeLLM, that is $E[p_{ij}|\theta_{i},\delta_{j},a_{j}]$, through their cumulative distributions across tasks. Besides observations of the cumulative distributions, we apply the Anderson-Darling (AD) (Scholz and Stephens, 1987) test [^2] The Anderson-Darling (AD) test is a statistical test to determine whether two independent samples come from the same (but unspecified) continuous distribution. The result of the test is a “p-value” which represents the probability of observing differences among the sample distributions as extreme as those measured, assuming both samples come from the same underlying distribution. A lower p-value suggests stronger evidence that the samples come from a different distribution. Generally, a threshold of 0.05 (95%) is selected., a statistical test to assess the difference between the two empirical distributions, between the distributions across benchmarks for the same CodeLLM. Then, using the obtained discriminant $a_{j}$ and difficulty $\delta_{j}$, as calculated by the IRT model, we will draw a portrait, task-wise, of both benchmarks. This will allow us to describe tasks present in the benchmark qualitatively.
RQ2: Do tasks’ topics affect their characteristics? In this RQ, we explore how tasks’ topics can affect how the difficulty/discriminant parameters are evaluated. We chose to explore at the topic level as it gives a good trade-off between a more nuanced analysis than the benchmark level while allowing us, compared to individual tasks analysis, to draw some conclusions over a group of tasks. To obtain meaningful topics, we will leverage the topic modelling approach with BERTopic (Grootendorst, 2022), as was used in similar studies when dealing with natural language data (Hämäläinen et al., 2023; Fütterer et al., 2023; Shen et al., 2023). BERTopic is a topic modelling technique that uses transformer-based embeddings (like BERT) to capture the semantic meaning of documents. It first converts documents into dense vectors, then reduces their dimensionality using UMAP (McInnes et al., 2018) to make clustering feasible. These vectors are clustered using HDBSCAN (McInnes et al., 2017), which groups similar documents without requiring the number of topics in advance. Each cluster’s most representative words are extracted using a class-based TF-IDF (c-TF-IDF) method to form interpretable topics. The result is a set of meaningful topics with associated keywords and document groupings.
@@
For each benchmark, the topic modelling clusters all tasks using the prompts of Level 1 before rephrasing. We use this specific prompt type to have similar formatting across tasks while being as close as possible to the original prompt, as mentioned in Section 3.2. We consider a topic to be valid only if it contains at least three tasks. This criterion ensures that in ClassEval, we avoid ending up with an excessive number of topics, each containing only two tasks, for example, which could occur due to the sampling methodology used to select tasks from the benchmark. As the process will inherently label some tasks as “noise”, noisy tasks are not considered in the analysis. This results in 17 topics for HumanEval+ and 21 topics for ClassEval. Those topics constitute tasks that have a similar functionality/goal. For the topics’ names, we used a refined version of the names provided by BERTopic. To refine the name of the topic, the first two authors manually check the tasks assigned in each topic by BERTopic, notably the prompt given as a description of the tasks. Based on this information, the name was updated by the authors independently before meeting to agree on a common name. The goal of this step was to give a topic name both more explicit than what BERTopic will assign as well as more general. For instance, one of ClassEval’s original topic names is “sql_select_female_under_age_table_name_query”. This is a BERTopic aggregation of keywords identified in functions’ descriptions dealing with SQL. To make it more understandable, and since no other topic focused on SQL requests, it was decided after independent analysis and a discussion to modify the name to “SQL Requests”, which is more understandable. Then, for each topic, the mean accuracy of each CodeLLM as well as the difficulty and discriminant among all tasks in the topic is calculated and analysed.
@@
RQ3: Do programming constructs affect tasks’ characteristics? In this RQ, we aim to explore the relationship between the difficulty or discriminative power of a task for a CodeLLM and the program constructs (e.g., If statements, For-loops, Function Calls, etc.) present in the generated code snippets used to solve the task. To achieve this, we will analyse the Abstract Syntax Tree (AST) of each code fragment generated for the task. Selecting only correct code samples (i.e. the ones passing all the tests) would be limiting, as code samples can be wrong to different degrees depending on the bugs affecting the code sample (Tambon et al., 2024). It could also bias the analysis towards code fragments obtained for higher-level prompts that are more likely to generate correct code. However, we also should not take any code fragment, as we may end up with generated code not even try to solve the task if the CodeLLM misunderstood the prompt. To strike a balance, we will only consider code fragments that are similar enough to a correct one. We do so by using CodeBLEU (Ren et al., 2020), which measures the syntactical correctness of code samples. We thus measure the similarity between each code fragment and a correct code fragment that passed the tests and only retain the code fragments with a similarity greater than a pre-defined threshold. Note that there is always one correct code fragment per task as a reference, which is provided in the benchmark. We chose CodeBLEU as it is a widely used metric in code generation tasks (Lu et al., 2021; Mastropaolo et al., 2023; Lozhkov et al., 2024; Wang et al., 2021). Once the code fragments that are similar enough have been obtained, we can proceed to collect the AST node type. We used the ast package from the Python standard library. For each node extracted in that way, we register the node type as per the AST. Once the nodes are extracted, as the number of lines of code could bias the results (e.g., harder tasks may require more lines of code to be solved), we normalise by the number of lines of code of each fragment. Finally, we calculate the average number of each program construct across all code fragments within the same task. To prevent averaging over too few code fragments, we only retain the tasks for which all models produced at least 10% of similar enough code fragments. Similarly, to prevent calculating correlations for node types that are barely represented across tasks, we ensure that at least 10% of the tasks include a particular node type to include it for comparison. This gives us the number of each node type, across tasks and CodeLLMs, which will be compared with the difficulty/discriminant of each task using the Kendall-$\tau$ (Kendall, 1938) [^3] Kendall’s tau test is a non-parametric test that assesses the strength and direction of association between two ranked (ordinal) variables. The tau value measures the correlation between the rankings, ranging from –1 (perfect inverse agreement) to +1 (perfect agreement), with 0 indicating no association. A p-value can be calculated as done in the AD test. In that case, a lower p-value provides stronger evidence that there is a statistically significant association between the two variables. to measure the correlation while handling ties. The node types with a significantly positive correlation will signify an increasing trend with difficulty. Furthermore, we will also use the AD test to check whether the distributions of the top/bottom 50% difficult/discriminant tasks exhibit different behaviour for each programming construct and each CodeLLM.
The threshold to use in CodeBLEU is a critical point, as it can drastically affect the outcome. Indeed, a higher threshold will mean that code fragments are likely very close to a correct solution, but their number will be low, which might limit any test relevance. On the other hand, a low threshold means a sufficient number of code fragments per task, but an increased possibility of including code fragments not implementing the desired functionality. As a manual analysis to establish a reliable threshold when dealing with such a large number of samples is impractical, we will instead choose to cross-compare using multiple thresholds. That is, we will retain code fragments with a similarity above 0.4, 0.5, 0.6, 0.7, 0.8, respectively. Bounds were chosen empirically: 0.4 is the threshold for which CodeBLEU authors established a correlation with an average human judgment of 4 out of 5, and a similarity above 0.8 (e.g. 0.9) would return too few samples. Then, when calculating the Kendall-$\tau$ and the AD test for each programming construct, we will only retain those programming constructs for which there is a statistical significance across all thresholds at the 0.05 threshold. We employ the Holm-Bonferroni (Holm, 1979) correction to ensure that a significant p-value is observed across all thresholds, that is, we iteratively verify $p_{i}<0.05/(5-(i-1))$ for i in [1; 5] and $p_{i}$ is the most significant p-value at iteration i. Doing so, we ensure that the observed statistical significance is robust to the choice of the threshold. In our experiment, this process resulted in between 80% and 94% for ClassEval and between 88% and 99% for HumanEval+ of tasks being considered in our analysis, depending on the threshold used. Moreover, while very high difficulty tasks (¿0.9) were generally removed because of too low a number of valid fragments, we ensured that at least 15% of tasks with a difficulty of 0.5 or more were present, no matter the CodeBLEU threshold used.
@@
RQ4: Do human annotators and CodeLLMs rate tasks’ difficulty similarly? In our final RQ, we aim to compare task difficulty assessments made by human annotators with those generated by our approach. To achieve this, we collect human-labelled difficulty ratings for both HumanEval+ and ClassEval by employing paid developers as crowd workers on the Prolific platform [^4] https://www.prolific.com/, which was used in previous studies (Zid et al., 2024). To have the crowd-workers estimate the difficulty, the most straightforward approach would be to have them implement the selected tasks. However, a pilot study with 10 developers for 12 tasks we conducted revealed that several participants are inclined to use LLMs to code for themselves, despite it being forbidden as advertised in the study. Given our limited control over how recruited participants might use LLMs, we adopted an alternative experimental setup that leverages effort estimation techniques, such as Planning Poker (Grenning, 2002), which is familiar to developers. Since our study involves a crowd of workers, we followed the experimental design proposed by Alhamed et al. (Alhamed and Storer, 2021), which focuses on effort estimation when relying on a crowd of non-experts. Their approach simulates real-world effort estimation processes but eliminates the need for developers to resolve estimation conflicts. The method is structured into iterative rounds. In the initial round, individual participants are asked to estimate the difficulty of a task through a time-based scale (e.g., one hour or one day) using solely the information included in the task description. Participants are also required to provide comments justifying their estimates. Once a sufficient number of responses are collected, the level of agreement among participants is calculated. If the agreement exceeds a predefined threshold, the task’s effort estimation is determined by taking the median of the responses. If the threshold is not met, a subsequent round is conducted with a new set of participants. These participants are provided with the task information, along with the comments and results from the previous round. This process repeats until either an agreement is reached or a maximum number of rounds is completed.
@@
In this study, we implemented Alhamed et al.’s approach with $5$ participants per task for a given round and a threshold of agreement level set to $0.6$ (i.e., Substantial Agreement). Given that we have a limited budget (as crowd-workers are paid based on the time spent), we aimed for $50$ developers in the first round. This results in $120$ tasks (out of the $364$ of both benchmarks) being considered, which were sampled at random. Tasks were then sorted into $10$ questionnaires of $12$ assessments ($6$ for HumanEval+ and $6$ for ClassEval). To balance the time required to complete the questionnaires, we categorised tasks into three groups based on the number of lines of code and the cyclomatic complexity of their oracles before sampling. All questionnaires can be found in our replication package (rep, 2024). Each assessment is decomposed into several parts: First, several prompts for the same tasks are provided to the participants, who have to summarise what the task is about. This allows us to check that the participant properly read the task. Prompts given to the participants, for each task, were sampled from the prompts CodeLLMs were evaluated on. Thus, both humans and LLMs had access to a similar description of the task. We made sure to always include the original prompt in the sampled prompts given to participants. Then, the participant is invited to estimate the duration it would take them (from 1 minute to over 40 minutes) as well as the subjective difficulty of the task (using a Likert scale from ”Very Easy” to ”Very Difficult”). They then had to explain their choice, based on considerations such as the time it would take to look for a particular concept and how to use a given library. At the beginning of the questionnaire, participants were given two examples to clarify expectations. Responses were collected until $5$ participants provided acceptable answers. To ensure quality, we manually reviewed responses, verifying that participants provided meaningful explanations (e.g., avoiding one-word answers for reasoning questions or random difficulty/time estimations). In particular, while we rely on the time estimation for our analysis, we still ask for the subjective difficulty. Analysing the trend between the two variables for each participant might help in potentially detecting incoherence (e.g., Easy tasks with 40 minutes estimation and Hard tasks with 10 minutes), which could signal random filling of the questionnaire. We followed Alhamed et al. (Alhamed and Storer, 2021) procedure until all tasks had been assessed with substantial agreement. We then analysed the difference between human-assessed difficulty and the difficulty for CodeLLMs obtained using our IRT-based approach. Since human assessments are discrete, we discretised the difficulty scores obtained via IRT within the interval [0,1] for comparison.
Applying the above process resulted in a total of 72 participants. We used prolific filters to only consider developers with at least 3 years of experience. During the questionnaire, the participants were further asked to report their experience using three categories: ¡ 5 years, 5 - 10 years, ¿ 10 years. The experience of the whole pool of the participants is as follows: 33,33% (24) of the participants have ¡ 5 years of experience, 44,44% (32) have between 5 and 10 years and 22,22% (16) have over 10 years.